A consultant traveled 8 hours to attend a meeting. The return trip took only 7 hours because the speed was 3 miles per hour faster. What was the consultant’s speed each way
Answer by stanbon(48551)

A consultant traveled 8 hours to attend a meeting. The return trip took only 7 hours because the speed was 3 miles per hour faster. What was the consultant’s speed each way
---------------------
To the meeting DATA:
Time = 8 hrs ; rate = x mph ; distance = 8x miles
-------------------
From the meeting DATA:
Time = 7 hrs ; rate = x+3 mph ; distance = 7(x+3) miles
-------------------------------------
EQUATION:
distance to = distance from
8x = 7x+21
x = 21 mph (rate to the meeting)
x+3 = 24 mph ( rate from the meeting)
===============================
Cheers,
Stan H.